# Resolving Interchange Reconnection Dynamics in a Fan-Spine-like Topology Observed by Solar Orbiter

### v1

@@
Interchange reconnection is believed to play a significant role in the production of solar jets and solar wind. However, the dynamics of interchange reconnection in the low corona might be more complex than recognized before in higher temporal and spatial resolutions. Using unprecedentedly high-resolution observations from the Extreme Ultraviolet Imager (EUI) onboard the Solar Orbiter, we analyze the dynamics of interchange reconnection in a small-scale fan-spine-like topology. Interchange reconnection that continuously occurs around the multi-null points of the fan-spine-like system exhibits a quasi-periodicity of 200 s, nearly covering the entire evolution of this system. Continuous evolution and reversal of multiple current sheets are observed over time near the null point. These results reveal that the dynamics of interchange reconnection are likely modulated by the emerging magnetic structures, such as mini-filaments and emerging arcades. Moreover, a curtain-like feature with a width of 1.7 Mm is also observed near the interchange reconnection region and persistently generates outflows, which is similar to the separatrix curtain reported in the pseudo-streamer structure. This study not only demonstrates the complex and variable reconnection dynamics of interchange reconnection within small-scale fan-spine topology but also provides insights into the self-similarity of magnetic field configurations across multiple temporal and spatial scales.
@@
Interchange reconnection represents a fundamental physical process in the solar atmosphere. It is characterized by the reconnection between closed magnetic field lines and adjacent open magnetic fields of opposite polarity (Fisk et al., 1998; Crooker et al., 2002). It can lead to a rapid conversion of magnetic energy, enabling the transfer of mass and energy along magnetic field lines. Meanwhile, interchange reconnection can cause plasma heating and particle acceleration, allowing energetic electrons to escape into interplanetary space along open magnetic field lines. Additionally, interchange reconnection is also an effective way to release the magnetic twist from closed magnetic fields into the interplanetary space environment (e.g., Bale et al., 2019; Sterling & Moore, 2020; Sterling et al., 2020). In the solar corona, diverse magnetic field configurations are conducive to the occurrence of interchange reconnection. Therefore, characterizing interchange reconnection dynamics in different magnetic field configurations is crucial for the triggering mechanisms for solar jets (e.g., Shibata et al., 1992; Krucker et al., 2011), as well as understand the origin and properties of the solar wind (e.g., Masson et al., 2014; Pellegrin-Frachon et al., 2023).
@@
Pseudo-streamers, as one of the large-scale coronal structures, provide a favorable environment for the occurrence of interchange reconnection. Pseudo-streamer separates open field lines of the same polarity and contains two closed magnetic flux below its cusp (Wang et al., 2007), displaying a -type structure. The pseudo-streamers are usually located more than 1.5 solar radii above the solar surface and are formed by multiple null points linked by separators (e.g., Titov et al., 2011, 2012; Seaton et al., 2013; Masson et al., 2014). In such a configuration, a strong gradient of magnetic connectivity exists among open and closed magnetic fields (e.g., Masson et al., 2014; Scott et al., 2021; Pellegrin-Frachon et al., 2023), known as Separatrix-Web (S-Web) and predicted to favor the occurrence of interchange reconnection. The S-Web model was proposed by Antiochos et al. (2011), Titov et al. (2011), and Linker et al. (2011), showing a network of separatrix surfaces and quasi-separatrix layers (QSLs). Simulation studies demonstrate that a vertical fan surface emanates from the null points, referred to as a separatrix curtain, which exhibits a curtain-like shape (e.g., Masson et al., 2014; Scott et al., 2021; Pontin & Priest, 2022). The separatrix curtain is surrounded by QSLs, which results from a rapid divergence of the field lines in the vicinity of the null points associated with the separatrix surfaces (Titov et al., 2011). The interchange reconnection occurring in pseudo-streamers is a gradual physical process (Masson et al., 2014) that can last for several hours without any drastic changes (e.g., Seaton et al., 2013). The pseudo-streamer topology, which may more accurately reflect the actual magnetic structure of the system, is thought to play a significant role in the generation of the slow solar wind. Therefore, understanding the dynamics of this system is crucial for analyzing the mass and energy release of the slow solar wind. Recently, Chitta et al. (2023) utilized middle coronal observations to investigate the dynamics of pseudo-streamer systems, providing observational evidence that the reconnection dynamics along the S-Web are responsible for the release of slow solar wind plasma. However, observing such a dynamic process of the pseudo-streamer system in high coronal still remains challenging.
@@
The fan-spine structure is defined by a dome-shaped fan separatrix surface, a magnetic null point positioned at its apex, and the inner and outer spines passing through the null point (e.g., Priest & Forbes, 2002; Li et al., 2021; Pontin & Priest, 2022). Similar to the large-scale pseudo-streamers, a smaller-scale fan-spine structure is also a desirable place for the occurrence of interchange reconnection. Recent observations revealed that the magnetic null point position of fan-spine configurations is approximately 16 Mm above the solar surface (e.g., Cheng et al., 2023; Duan et al., 2024b). The fan-spine topology is thus associated with various small-scale solar phenomena, such as solar jets (e.g., Liu et al., 2011; Wang & Liu, 2012; Hong et al., 2017; Shen et al., 2019; Duan et al., 2022), solar flares (e.g., Sun et al., 2013; Li et al., 2018; Song & Tian, 2018; Zhong et al., 2019; Yang et al., 2020b), ultraviolet bursts (e.g., Chitta et al., 2017; Chen et al., 2019b; Bhatnagar et al., 2025b), Ellerman bombs (e.g., Bhatnagar et al., 2025a), macrospicules (e.g., Duan et al., 2023), small-scale spiral jets (e.g., Cheng et al., 2023; Li et al., 2025b), jetlets (e.g., Kumar et al., 2022; Yang et al., 2025), and coronal bright points (e.g., Zhang et al., 2012; Færder et al., 2024; Nóbrega-Siverio et al., 2025). Currently, an increasing number of observations and simulations support that jets and flares could result from magnetic reconnection between the magnetic flux rope (or filament) initially confined below the fan and the magnetic field of the outer fan (e.g., Jiang et al., 2013; Song et al., 2018; Wyper et al., 2017, 2018; Yang et al., 2020a; Zhou et al., 2021; Duan et al., 2024b; Yang et al., 2024a). After the eruption, the remains of an inner bright patch encircled by a circular flare ribbon can be observed on the solar surface (e.g., Masson et al., 2009; Liu et al., 2015; Hou et al., 2019; Ning et al., 2022; Zhang, 2024), and these features indicate the footpoint positions of the inner spine and the dome-like fan separatrix surface, respectively. The lifetime of circular ribbon flares ranges from 4 to 205 minutes, with a mean value of approximately 50 minutes (Zhang et al., 2022). The speeds and ejection heights of solar jets can reach several hundred kilometers per second and several hundred Mm, respectively.
@@
Interchange reconnection around the null point in the fan-spine system is also an active field of theoretical and numerical research, which aids in understanding the transfer of magnetic flux across different topological regions (e.g., Pariat et al., 2009; Masson et al., 2012; Fuentes-Fernández & Parnell, 2013; Pontin et al., 2013; Pontin & Priest, 2022). Traditional studies have generally considered that fan-spine structures contain a single null point (e.g., Pariat et al., 2009; Wyper et al., 2017), which differs from pseudo-streamer structures that are formed by several null points linked by separators (e.g., Titov et al., 2012). However, research by Duan et al. (2024b) reported that fan-spine topology may contain more than one null point (potentially multiple null points, or QSLs, or separators). Duan et al. (2024b) found that the position of a current sheet is along the QSL connecting two null points. Moreover, the small-scale evolving jets can also exhibit various dynamic processes at their base, such as multiple small-scale eruptions, plasma flows, and moving blobs (e.g., Mandal et al., 2022; Long et al., 2023). Thus, the dynamics of interchange reconnection within small-scale fan-spine structures may exhibit more complex reconnection behaviors, necessitating further investigations with higher resolution.
@@
In this paper, we utilize high temporal and spatial resolution data acquired by the Extreme Ultraviolet Imager (EUI; Rochus et al. (2020)) on the Solar Orbiter (SO; Müller et al. (2020)) to investigate the dynamics of interchange reconnection that continuously occurs around the null point of a fan-spine-like structure. Moreover, we also identify a curtain-like feature similar to the large-scale pseudo-streamer within a small-scale fan-spine-like structure.
@@ 2 Observations
On 2024 April 5, Solar Orbiter (SO) was positioned approximately 0.29 AU from the Sun, at an angle of 83 degrees between SO and Earth relative to the Sun (e.g., Tan et al., 2025). The event of interest occurred at the western limb of the Sun as observed by SO, rendering it completely invisible from Earth. Consequently, all data utilized in this study were obtained from SO. The Extreme Ultraviolet Imager (EUI) aboard SO consists of three telescopes: the Full Sun Imager (FSI) and two High-Resolution Imagers (HRIEUV and HRILyα). HRIEUV 174 Å images provide the dynamic evolution of the fan-spine-like structure and the complex interchange reconnection process around the null point within the fan-spine-like system. The HRIEUV 174 Å observations feature a time cadence of 16 s, and a pixel scale of 0.″49, with an effective exposure time of 2 s. A pixel corresponds to 0.108 × 0.108 Mm2 on the Sun. The imaging observations in two channels (FSI 174 Å and FSI 304 Å ) from the FSI also enable us to study the plasma motions from a larger field of view. The temporal cadence and spatial pixel scale of FSI data are 10 minutes and 4.44″, respectively. We used cross-correlation for data alignment and jitter elimination.
@@ 3 Results
Figure 1 (a) and (b) present an overview of a fan-spine-like structure observed at the solar limb from the perspective of Solar Orbiter. A closer view of this structure, indicated by the cyan box in Figure 1 (b), is shown in Figure 1 (c) after a 115° counterclockwise rotation to achieve an upright orientation. Figure 1 (c) reveals that the scale of the fan dome is approximately 15 Mm at the solar limb. At this time, plasma outflows along the outer spine of the fan-spine system display the morphology of an anemone solar jet, characterized by an inverted-Y shape comprising a bright, straight jet spire and a jet base. This inverted-Y feature is thought to be the consequence of reconnection between small-scale magnetic arcades and the unipolar background fields (e.g., Tian et al., 2014a, 2018; Tang et al., 2025). Given the location of this structure at the solar limb and the associated projection effects, we did not use PHI (Solanki et al., 2020) for magnetic field extrapolation (e.g., Petrova et al., 2024) to obtain the fan-spine feature. However, in the context of magnetic reconnection in three-dimensional geometries, the simplest anticipated outcome of flux emergence into a locally unipolar region is the formation of a fan-spine magnetic field configuration (e.g., Török et al., 2009). Thus, the fan-spine topology can represent the 3D magnetic structure of the straight anemone jet (e.g., Shen, 2021).
@@
Figure 2 exhibits the evolution of the fan-spine-like topology, as observed in HRIEUV 174Å images. At approximately 22:00:44 UT, a small, dark, closed-loop system emerges to the right of the coronal plume region. This dark loop could represent an arch filament system carrying cool material (e.g., Bruzek, 1967; Chen et al., 2018; Su et al., 2018). By 22:09:48 UT, the emerged loop gradually expands toward the background coronal loop region. The speed at which the magnetic loop carrying cold material approached the background field is approximately 2 km s-1, as shown in Figure 2 (f). At 22:13:48 UT, an X-point, the signal of interaction reconnection between two systems at the reconnection site, is marked by the yellow arrows. This interaction results in the formation of a new closed loop (denoted by the white arrow), as shown in Figure 2 (d). Afterward, a weak sheet-like feature develops near the X-point, connecting the new loop and the open field, as denoted by the cyan arrow in Figure 2 (d). Such a sheet-like structure delineates the current sheet associated with the reconnection configuration, as suggested by previous observational studies (e.g., Tian et al., 2014b; Xue et al., 2016; Shen et al., 2017; Chen et al., 2019a; Hou et al., 2021; Lu et al., 2022; Ding & Zhang, 2024; Duan et al., 2024b). Moreover, bright plasma blobs can also be observed within the current sheet, as indicated by the pink arrows in Figure (d), as reported by previous studies (e.g, Li et al., 2016; Yan et al., 2022; Hou et al., 2024; Nóbrega-Siverio et al., 2025; Tang et al., 2025). The plasma blobs may be generated during the interchange reconnection as a result of tearing instability, as indicated by the simulation studies conducted by Bhattacharjee et al. (2009). As the interaction becomes more dynamic, a common footpoint region shared by the emerged loop and the newly formed loops is observed, collectively displaying a fan-spine-like morphology, as outlined by dashed lines in Figure 2 (e). The interaction between this newly emerging magnetic flux and the adjacent background field evolves into a fan-spine morphology, consistent with the previous work of Török et al. (2009) and Duan et al. (2024b). In this event, due to the position of our interested event at the solar western limb, lacking the magnetic field data, we cannot employ magnetic field extrapolation. Thus, we refer to this event as the fan-spine-like structure.
@@
Figure 3 presents the observational dynamic features associated with the transition from current sheet 2 (CS2) to current sheet 3 (CS3). At about 22:17:16, a bright emission appears beneath filament 1, located at the side of the fan-spine-like structure. Meanwhile, a CS2 feature appears within the fan-spine-like structure. Although the EUI image only provides the projected trajectory of filament 1 and CS2, it is clear that they move upward together (see the animation). The upward movement of the CS2 could be a reflection of larger-scale magnetic field relaxation in that region.Ultimately, filament 1 erupts, as evidenced by its structural changes from a closed magnetic flux system to an open field, as displayed in Figure 3 (c). Another filament (filament 2) is observed at 22:21:32 UT. About two minutes later, filament 2 is activated and erupts in the direction of CS2. This result may be triggered by CS2, effectively eliminating the restraining magnetic arcade overlaying the underlying filament 2. At 22:24:12 UT, a bright linear feature, indicated by the white arrow, forms a connection between the base of the jet spire and the top of the post-flare arcade. This feature is similar to the flare current sheet formed in the wake of the filament eruption(Antiochos et al., 1999; Wyper et al., 2017), referred to as the CS3. Figure 3 (d1)(f1) exhibits the schematic diagrams corresponding to panels (d)(f) and the possible connectivity of filament 2.
@@
By constructing a time–distance diagram along the arrow S3 (see Figure 3 (f)), it can be observed continuous occurrence of plasma outflows and a significant jet eruption ( 22:23 UT), as displayed in Figure 3 (i3). We also employed the wavelet and Fourier analysis method (Torrence & Compo, 1998; Auchère et al., 2016) to obtain the periods of the continuous plasma outflows. We extracted a light curve from S3, with the location marked by the red dashed line in Figure 3 (i3). The result is illustrated in Figures 7 (a1)(a3). The speeds of these plasma outflows are presented in the time-distance diagram along the ejection to the highest position, as exhibited in Figures 6 (a) and (d).
@@
Figure 3 (i1) and (i2) show the time–distance diagrams along cyan arrows S1 and S2 in panel (f). At around 22:23 UT, it can be clearly seen that CS2 transforms into a two-loop system moving in different directions (see the animation), while CS3 appears. It is worth noting that at this moment, filament 2 erupts toward CS2, as indicated by the red arrow in Figure 3 (i1). CS2 is observed from approximately 22:17 UT to 22:23 UT, during which its width ranged from 0.6 to 0.8 Mm, consistent with prior statistical analyses (Ding & Zhang, 2024; Ding et al., 2024). The length of CS2 seems to grow over time, from an initial 3 Mm to a final 5 Mm before disappearing. According to the Sweet-Parker mechanism (Priest & Forbes, 2000), the reconnection rate (the inflow Alfvn Mach number (MA)) is theoretically proportional to the ratio of the current sheet thickness (d) to its length (L) (e.g., Xue et al., 2016; Chen et al., 2019a; Yang et al., 2024b). Based on these measurements, the reconnection rate of the CS2 is estimated to be in the range of 0.12 to 0.2. During the period from 22:23 to 22:40 UT, CS3 displayed a width of approximately 0.4 Mm to 0.6 Mm, while its length varied from 2.7 Mm to 7.4 Mm. Unlike CS2, CS3 exhibited a contraction phase following its initial elongation. The estimated reconnection rate of CS3 ranged from 0.12 to 0.17. Both current sheets show an upward movement trend, with CS2 and CS3 moving at speeds of 5.5 km/s and 1.5 km/s, respectively.
@@
Figure 4 illustrates the dynamic evolution of the interchange reconnection in the outflow region of CS3. A set of magnetic loops is clearly visible to the left of the jet spire, as shown in Figure 4(a). These loops are reconnected loops formed in the outflow region of CS2. At 22:25:16 UT, this set of magnetic loops clearly exhibits three distinct groups of field lines, indicated by the yellow arrows. Between 22:24:28 UT and 22:25:32 UT, these loops move toward the jet’s spire, exhibiting a slipping motion. They reconnect with the surrounding open field lines at CS3, with the reconnection point indicated by the purple arrow in panel (d). At 22:26:36, the newly formed loops after interchange reconnection are shown by the cyan arrow. Eventually, the closed field lines on the left of the jet spire change into open field lines through the dynamic evolution of interchange reconnection, revealing a curtain-like feature (see panels (f)(i)). The width of this curtain-like feature is approximately 1.7 Mm. By constructing a time-distance diagram along the curtain-like structure from arrows S4S6, we can observe the continuous plasma outflows along the curtain feature. The average velocity of these plasma outflows is 80 km s-1. We extracted the light curves from S4, S5, and S6, with the location marked by the red dashed lines in Figures 4 (i1)(i3). The results of wavelet and Fourier analysis are shown in Figures 7 (b)(d).
@@
Figure 5 shows the ultimate breakdown process of this fan-spine-like system. CS3 gradually approaches the post-flare loop, as shown in Figure 5 (a)(c). This may result from the migration of the null point in the fan-spine system, caused by changes in gas pressure and magnetic pressure. Subsequently, another filament, referred to as filament 3, appears below the post-flare loop, as displayed in Figure 5 (d). Together, they exhibit dynamic evolutionary characteristics of uplift and interact with each other (see Figure 5 (e)). Eventually, the post-flare loop and filament 3 transition from closed magnetic flux to open flux at 22:47:40 UT.

### v2

@@
Interchange reconnection is thought to play a significant role in the production of solar jets and the solar wind. The dynamics of interchange reconnection in the low corona might be more complex than recognized before in higher temporal and spatial resolutions, however. Using unprecedentedly high-resolution observations from the Extreme Ultraviolet Imager (EUI) on board the Solar Orbiter, we analyzed the dynamics of interchange reconnection in a small-scale fan-spine-like topology. The interchange reconnection that continuously occurs around the multinull points of the fan-spine-like system exhibits a quasi-periodicity of 200 s that nearly covers the entire evolution of this system. Continuous evolution and a reversal of multiple current sheets are observed over time near the null point. These results reveal that the dynamics of interchange reconnection is likely modulated by emerging magnetic structures, such as mini-filaments and emerging arcades. Moreover, a curtain-like feature with a width of 1.7 Mm is also observed near the interchange reconnection region and persistently generates outflows. This is similar to the separatrix curtain reported in the pseudostreamer structure. This study not only demonstrates the complex and variable reconnection dynamics of interchange reconnection within a small-scale fan-spine topology, but also provides insights into the self-similarity of magnetic field configurations across multiple temporal and spatial scales.
@@
Interchange reconnection represents a fundamental physical process in the solar atmosphere. It is characterized by the reconnection between closed magnetic field lines and adjacent open magnetic fields of opposite polarity (Fisk et al., 1998; Crooker et al., 2002). It can lead to a rapid conversion of magnetic energy that enables the transfer of mass and energy along magnetic field lines. Meanwhile, interchange reconnection can cause plasma heating and particle acceleration, which allows the escape of energetic electrons into interplanetary space along open magnetic field lines. Additionally, interchange reconnection is also an effective way to release the magnetic twist from closed magnetic fields into interplanetary space (e.g., Bale et al., 2019; Sterling & Moore, 2020; Sterling et al., 2020). In the solar corona, diverse magnetic field configurations are conducive to the occurrence of interchange reconnection. It is therefore crucial to characterize the interchange reconnection dynamics in different magnetic field configurations for understanding the triggering mechanisms for solar jets (e.g., Shibata et al., 1992; Krucker et al., 2011) and the origin and properties of the solar wind (e.g., Masson et al., 2014; Pellegrin-Frachon et al., 2023).
@@
Pseudostreamers are large-scale coronal structures and provide a favorable environment for the occurrence of interchange reconnection. Pseudostreamers separate open field lines of the same polarity and contain two closed magnetic fluxes below their cusp (Wang et al., 2007) with a -type structure. The pseudostreamers are usually located more than 1.5 solar radii above the solar surface and are formed by multiple null points linked by separators (e.g., Titov et al., 2011, 2012; Seaton et al., 2013; Masson et al., 2014). In this configuration, the gradient of magnetic connectivity among open and closed magnetic fields is strong (e.g., Masson et al., 2014; Scott et al., 2021; Pellegrin-Frachon et al., 2023). This is known as a separatrix-web (S-Web) and is predicted to favor the occurrence of interchange reconnection. The S-Web model was proposed by Antiochos et al. (2011), Titov et al. (2011), and Linker et al. (2011), who showed a network of separatrix surfaces and quasi-separatrix layers (QSLs). Simulation studies demonstrated that a vertical fan surface emanates from the null points. This is referred to as a separatrix curtain because it has a curtain-like shape (e.g., Masson et al., 2014; Scott et al., 2021; Pontin & Priest, 2022). The separatrix curtain is surrounded by QSLs, which results from the rapid divergence of the field lines in the vicinity of the null points associated with the separatrix surfaces (Titov et al., 2011). The interchange reconnection that occurs in pseudostreamers is a gradual physical process (Masson et al., 2014) that can last for several hours without any drastic changes (e.g., Seaton et al., 2013). The pseudostreamer topology, which probably more accurately reflects the actual magnetic structure of the system, is thought to play a significant role in the generation of the slow solar wind. It is therefore crucial to understand the dynamics of this system for the analysis of the mass and energy release of the slow solar wind. Recently, Chitta et al. (2023) used middle coronal observations to investigate the dynamics of pseudostreamer systems, providing observational evidence that the reconnection dynamics along the S-Web are responsible for the release of slow solar wind plasma. However, it remains challenging to observe this dynamic process of the pseudostreamer system in the high corona.
@@
The fan-spine structure is defined by a dome-shaped fan separatrix surface, a magnetic null point positioned at its apex, and the inner and outer spines that pass through the null point (e.g., Priest & Forbes, 2002; Li et al., 2021; Pontin & Priest, 2022). Similar to the large-scale pseudostreamers, a smaller-scale fan-spine structure is also a favorable place for the occurrence of interchange reconnection. Recent observations revealed that the magnetic null-point position of fan-spine configurations is approximately 16 Mm above the solar surface (e.g., Cheng et al., 2023; Duan et al., 2024b). The fan-spine topology is thus associated with various small-scale solar phenomena, such as solar jets (e.g., Liu et al., 2011; Wang & Liu, 2012; Hong et al., 2017; Shen et al., 2019; Duan et al., 2022), solar flares (e.g., Sun et al., 2013; Li et al., 2018; Song & Tian, 2018; Zhong et al., 2019; Yang et al., 2020b), ultraviolet bursts (e.g., Chitta et al., 2017; Chen et al., 2019b; Bhatnagar et al., 2025b), Ellerman bombs (e.g., Bhatnagar et al., 2025a), macrospicules (e.g., Duan et al., 2023), small-scale spiral jets (e.g., Cheng et al., 2023; Li et al., 2025b), jetlets (e.g., Kumar et al., 2022; Yang et al., 2025), and coronal bright points (e.g., Zhang et al., 2012; Nóbrega-Siverio & Moreno-Insertis, 2022; Færder et al., 2024). An increasing number of observations and simulations currently support that jets and flares can result from magnetic reconnection between the magnetic flux rope (or filament) that is initially confined below the fan and the magnetic field of the outer fan (e.g., Jiang et al., 2013; Song et al., 2018; Wyper et al., 2017, 2018; Yang et al., 2020a; Zhou et al., 2021; Duan et al., 2024b; Yang et al., 2024a). After the eruption, the remains of an inner bright patch encircled by a circular flare ribbon are observed on the solar surface (e.g., Masson et al., 2009; Liu et al., 2015; Hou et al., 2019; Ning et al., 2022; Zhang, 2024), and these features indicate the footpoint positions of the inner spine and the dome-like fan separatrix surface. The lifetime of circular ribbon flares ranges from 4 to 205 minutes, with a mean value of approximately 50 minutes (Zhang et al., 2022). The speeds and ejection heights of solar jets can reach several hundred kilometers per second and several hundred megameters, respectively.
@@
Interchange reconnection around the null point in the fan-spine system is also an active field of theoretical and numerical research. It helps us to understand the transfer of magnetic flux across different topological regions (e.g., Pariat et al., 2009; Masson et al., 2012; Fuentes-Fernández & Parnell, 2013; Pontin et al., 2013; Pontin & Priest, 2022). Traditional studies have generally considered that fan-spine structures contain a single null point (e.g., Pariat et al., 2009; Wyper et al., 2017), which differs from pseudostreamer structures, which are formed by several null points linked by separators (e.g., Titov et al., 2012). However, Duan et al. (2024b) recently reported that the fan-spine topology probably contains more than one null point (potentially multiple null points, QSLs, or separators). Duan et al. (2024b) reported that the position of a current sheet is along the QSL that connects two null points. Moreover, small-scale evolving jets also exhibit various dynamic processes at their base, such as multiple small-scale eruptions, plasma flows, and moving blobs (e.g., Mandal et al., 2022; Long et al., 2023). Thus, the dynamic of interchange reconnection within small-scale fan-spine structures is probably more complex and requires further investigations with higher resolution.
@@
We used data with a high temporal and spatial resolution acquired by the Extreme Ultraviolet Imager (EUI; Rochus et al. 2020) on board Solar Orbiter (Müller et al. 2020) to investigate the dynamics of interchange reconnection that continuously occurs around the null point of a fan-spine-like structure. Moreover, we also identify a curtain-like feature similar to the large-scale pseudostreamer within a small-scale fan-spine-like structure.
@@ 2 Observations
On 2024 April 5, Solar Orbiter was positioned approximately 0.29 AU from the Sun at an angle of 83 degrees between Solar Orbiter and Earth relative to the Sun (e.g., Tan et al., 2025). The event of interest occurred at the western limb of the Sun as observed by Solar Orbiter, rendering it completely invisible from Earth. Consequently, all data we used were obtained from Solar Orbiter. The EUI on board Solar Orbiter consists of three telescopes: the Full Sun Imager (FSI), and two High-Resolution Imagers (HRIEUV and HRILy). HRIEUV 174 Å images provide the dynamic evolution of the fan-spine-like structure and the complex interchange reconnection process around the null point within the fan-spine-like system. The time cadence of HRIEUV 174 Å observations is 16 s, and the pixel scale is 0.″49, with an effective exposure time of 2 s. A pixel corresponds to 0.108 × 0.108 Mm2 on the Sun. The imaging observations in two channels (FSI 174 Å and FSI 304 Å ) from the FSI also enable us to study the plasma motions from a larger field of view. The temporal cadence and spatial pixel scale of FSI data are 10 minutes and 4.44″. We used a cross-correlation for the data alignment and jitter elimination.
@@ 3 Results
Figures 1(a) and 1(b) present an overview of a fan-spine-like structure observed at the solar limb from the perspective of Solar Orbiter. A closer view of this structure, indicated by the cyan box in Fig. 1(b), is shown in Fig. 1(c) after a 115° counterclockwise rotation to achieve an upright orientation. Figure 1(c) reveals that the scale of the fan dome is approximately 15 Mm at the solar limb. At this time, the morphology of the plasma outflows along the outer spine of the fan-spine system is that of an anemone solar jet, characterized by an inverted-Y shape comprising a bright, straight jet spire and a jet base. This inverted-Y feature is thought to be the consequence of reconnection between small-scale magnetic arcades and the unipolar background fields (e.g., Tian et al., 2014a, 2018). Given the location of this structure at the solar limb and the associated projection effects, we did not use PHI (Solanki et al., 2020) to extrapolate the magnetic field (e.g., Petrova et al., 2024) and obtain the fan-spine feature. In the context of magnetic reconnection in three-dimensional geometries, the simplest anticipated outcome of flux emergence into a locally unipolar region is the formation of a fan-spine magnetic field configuration (e.g., Török et al., 2009). Thus, the fan-spine topology represents the 3D magnetic structure of the straight anemone jet (e.g., Shen, 2021).
@@
Figure 2 shows the evolution of the fan-spine-like topology as observed in HRIEUV 174Å images. At approximately 22:00:44 UT, a small dark closed-loop system emerges to the right of the coronal plume region. This dark loop might represent an arch filament system carrying cool material (e.g., Bruzek, 1967; Chen et al., 2018; Su et al., 2018). By 22:09:48 UT, the emerged loop gradually expanded toward the background coronal loop region. The speed at which the magnetic loop carrying cold material approached the background field is approximately 2 km s-1, as shown in Fig. 2(f). At 22:13:48 UT, an X-point, the signal of interaction reconnection between two systems at the reconnection site, is marked by the yellow arrows. This interaction results in the formation of a new closed loop (denoted by the white arrow), as shown in Fig. 2(d). Afterward, a weak sheet-like feature developed near the X-point and connected the new loop and the open field, as denoted by the cyan arrow in Fig. 2(d). This sheet-like structure delineates the current sheet associated with the reconnection configuration, as suggested by previous observational studies (e.g., Tian et al., 2014b; Xue et al., 2016; Shen et al., 2017; Chen et al., 2019a; Hou et al., 2021; Lu et al., 2022; Ding & Zhang, 2024; Duan et al., 2024b). Moreover, bright plasma blobs are also observed within the current sheet, as indicated by the pink arrows in Fig. 2(d) and reported by previous studies (e.g, Li et al., 2016; Yan et al., 2022; Hou et al., 2024; Nóbrega-Siverio et al., 2025; Tang et al., 2025). The plasma blobs might be generated during the interchange reconnection as a result of a tearing instability, as indicated by the simulation studies conducted by Bhattacharjee et al. (2009). As the interaction becomes more dynamic, a common footpoint region shared by the emerged loop and the newly formed loops is observed that collectively displays a fan-spine-like morphology, as outlined by the dashed lines in Fig. 2(e). The interaction between this newly emerging magnetic flux and the adjacent background field evolves into a fan-spine morphology, consistent with Török et al. (2009) and Duan et al. (2024b). Because of the position of the event at the solar western limb, which lacks magnetic field data, we cannot extrapolate the magnetic field for this event. Thus, we refer to this event as the fan-spine-like structure.
@@
Figure 3 presents the observational dynamic features associated with the transition from current sheet 2 (CS2) to current sheet 3 (CS3). At about 22:17:16, a bright emission appears beneath filament 1, located at the side of the fan-spine-like structure, and the CS2 feature appears within the fan-spine-like structure. Although the EUI image only provides the projected trajectory of filament 1 and CS2, they clearly moved upward together (see the animation). The upward movement of the CS2 might be a reflection of larger-scale magnetic field relaxation in this region. Ultimately, filament 1 erupted, as evidenced by its structural changes from a closed magnetic flux system to an open field, shown in Fig. 3(c). Another filament (filament 2) is observed at 22:21:32 UT. About two minutes later, filament 2 was activated and erupted in the direction of CS2. This may be triggered by CS2 and effectively eliminates the restraining magnetic arcade overlaying the underlying filament 2. At 22:24:12 UT, a bright linear feature, indicated by the white arrow, connected the base of the jet spire and the top of the post-flare arcade. This feature is similar to the flare current sheet that formed in the wake of the filament eruption(Antiochos et al., 1999; Wyper et al., 2017) and is referred to as CS3. Figures 3(d1)3(f1) show the schematic diagrams corresponding to panels (d)(f) and the possible connectivity of filament 2.
@@
In the time–distance diagram along the arrow S3 (see Fig. 3(f)), plasma outflows occur continuously and a significant jet erupts ( 22:23 UT), as displayed in Fig. 3(i3). We also employed the wavelet and Fourier analysis method (Torrence & Compo, 1998; Auchère et al., 2016) to obtain the periods of the continuous plasma outflows. We extracted a light curve from S3 (the location is marked by the dashed red line in Fig. 3(i3)). The result is illustrated in Figs. 7(a1)7(a3). The speeds of these plasma outflows are presented in the time-distance diagram along the ejection to the highest position, as shown in Figs. 6 (a) and 6(d).
@@
Figures 3(i1) and 3(i2) show the time–distance diagrams along the cyan arrows S1 and S2 in panel (f). At around 22:23 UT, CS2 clearly transformed into a two-loop system moving in different directions (see the animation), while CS3 appeared. At this moment, filament 2 erupted toward CS2, as indicated by the red arrow in Fig. 3(i1). CS2 was observed from approximately 22:17 UT to 22:23 UT, during which its width ranged from 0.6 to 0.8 Mm, which is consistent with prior statistical analyses (Ding & Zhang, 2024; Ding et al., 2024). The length of CS2 appeared to grow over time from an initial 3 Mm to a final 5 Mm before it disappeared. According to the Sweet-Parker mechanism (Priest & Forbes, 2000), the reconnection rate (the inflow Alfvén Mach number (MA)) is theoretically proportional to the ratio of the current sheet thickness (d) to its length (L; e.g., Xue et al., 2016; Chen et al., 2019a; Yang et al., 2024b). Based on these measurements, the reconnection rate of the CS2 was estimated to be in the range of 0.12 to 0.2. During the period from 22:23 to 22:40 UT, the width of CS3 was approximately 0.4 Mm to 0.6 Mm, and its length varied from 2.7 Mm to 7.4 Mm. Unlike CS2, CS3 contracted for a time following its initial elongation. The estimated reconnection rate of CS3 ranged from 0.12 to 0.17. Both current sheets showed an upward movement trend, and CS2 and CS3 moved at speeds of 5.5 km/s and 1.5 km/s, respectively.
@@
Figure 4 illustrates the dynamic evolution of the interchange reconnection in the outflow region of CS3. A set of magnetic loops is clearly visible to the left of the jet spire, as shown in Fig. 4(a). These loops are reconnected loops that formed in the outflow region of CS2. At 22:25:16 UT, this set of magnetic loops clearly showed three distinct groups of field lines (indicated by the yellow arrows). Between 22:24:28 UT and 22:25:32 UT, these loops moved toward the jet spire in a slipping motion. They reconnected with the surrounding open field lines at CS3 (the reconnection point is indicated by the purple arrow in panel (d)). At 22:26:36, the newly formed loops after interchange reconnection are shown by the cyan arrow. The closed field lines on the left of the jet spire eventually changed into open field lines through the dynamic evolution of interchange reconnection in a curtain-like feature (see panels (f)(i)). The width of this curtain-like feature is approximately 1.7 Mm. The time-distance diagram along the curtain-like structure from arrows S4S6 shows the continuous plasma outflows along the curtain feature. The average velocity of these plasma outflows is 80 km s-1. We extracted the light curves from S4, S5, and S6 at the location marked by the dashed red lines in Figs. 4(i1)(i3). The results of the wavelet and Fourier analysis are shown in Figs. 7(b)7(d).
@@
Figure 5 shows the ultimate breakdown process of this fan-spine-like system. CS3 gradually approached the post-flare loop, as shown in Figs. 5(a)5(c). This might be caused by the migration of the null point in the fan-spine system as a result of changes in the gas pressure and magnetic pressure. Subsequently, another filament, referred to as filament 3, appeared below the post-flare loop, as displayed in Fig. 5(d). Together, they showed dynamic evolutionary characteristics of uplift and interacted with each other (see Fig. 5(e)). The post-flare loop and filament 3 eventually transitioned from a closed magnetic flux to an open flux at 22:47:40 UT.